# Representational Stability of Truth in Large Language Models

### v2

@@ Abstract
Large language models (LLMs) are widely used for factual tasks such as “What treats asthma?” or “What is the capital of Latvia?”. However, it remains unclear how stably LLMs encode distinctions between true, false, and neither-true-nor-false content in their internal probabilistic representations. We introduce representational stability as the robustness of an LLM’s veracity representations to perturbations in the operational definition of truth. We assess representational stability by (i) training a linear probe on an LLM’s activations to separate true from not-true statements and (ii) measuring how its learned decision boundary shifts under controlled label changes. Using activations from sixteen open-source models and three factual domains, we compare two types of neither statements. The first are fact-like assertions about entities we believe to be absent from any training data. We call these unfamiliar neither statements. The second are nonfactual claims drawn from well-known fictional contexts. We call these familiar neither statements. The unfamiliar statements induce the largest boundary shifts, producing up to $40\%$ flipped truth judgments in fragile domains (such as word definitions), while familiar fictional statements remain more coherently clustered and yield smaller changes ($\leq 8.2\%$). These results suggest that representational stability stems more from epistemic familiarity than from linguistic form. More broadly, our approach provides a diagnostic for auditing and training LLMs to preserve coherent truth assignments under semantic uncertainty, rather than optimizing for output accuracy alone.
Figure 1: Overview of Representational Stability Evaluation. A toy example demonstrating how we assess representational stability by (a) training a True vs. Not True probe on LLM activations with True (blue), False (green) and Neither (purple) veracity values and (b) retraining the probe with perturbed labels (i.e., redefining the operational definition of truth to include the Neither statements). We compare the similarity between the original (solid) and perturbed (dashed) decision boundaries and identify how many True statements flip to Not True after the perturbation (epistemic retractions) or, conversely, how many Not True statements flip to True (epistemic expansions). Stable veracity representations should have well-clustered activations that minimize the number of epistemic retractions and expansions.
@@ 1 Introduction
Large language models (LLMs) are increasingly used as sources of information, yet their behavior often blurs the line between knowledge and plausibility [alkhamissi2022review, turpin2023language, han-etal-2025-simple]. People expect experts to distinguish between True, False, and Neither statements, yet it remains unclear whether LLMs form similarly structured internal representations. The stability of internal veracity representations in LLMs, i.e., how consistently they encode truth and falsity across related statements, is crucial for reliability and safety [liu2023trustworthy, suzgun2025language, abbasi2024believe]. When this representational structure is unstable, LLMs often exhibit undesirable behaviors such as hallucinations [huang2025survey, han-etal-2025-simple].
LLMs can appear factually competent even when their internal veracity representations are weakly separated or inconsistent. Such instability is one explanation for why small prompt or context changes can affect an LLM’s answers [elazar2021measuring, li2025firm, abbasi2024believe], with recent work suggesting that epistemic familiarity also shapes an LLM’s confidence and self-evaluation [kadavath2022language]. Prior work follows two paths: representation-based probing, which examines whether true and false statements form separable clusters in activation space [burger2024truth, savcisens2025trilemma, marks2310geometry], and in-context analyses, which test how output varies under persuasion [wilie2024belief, xu-etal-2024-earth], phrasing [turpin2023language, lu2022fantastically], or jailbreak attacks [wei2023jailbroken]. However, we lack a unified approach for identifying which kinds of statements disrupt an LLM’s latent factual representation [harding2023operationalizing].
We address this gap by analyzing representational stability: the consistency of an LLM’s veracity representations under controlled perturbations of a probe’s training data (see Figure 1). Inspired by Leitgeb’s notion of $P$-stability [leitgeb2014stability], a property of belief systems requiring stability under small evidential changes, we treat a statement’s embedding as a belief state and use representation-based probes to identify truth directions.
We analyze three factual domains, City Locations, Medical Indications, and Word Definitions, and five statement types (True, False, Fictional, Synthetic, and Noise). Fictional and Synthetic statements represent distinct Neither cases. Fictional statements originate from familiar imaginary worlds likely present in training corpora, whereas Synthetic statements are automatically generated to ensure unfamiliarity. We train a probe on activations from sixteen open-source LLMs to learn a baseline True vs. Not True direction. For our probe, we use sAwMIL: a max-margin, multiple-instance probe designed to incorporate Neither statements [savcisens2025trilemma]. We then retrain the same probe under label perturbations (e.g., treating Fictional statements as True) to quantify shifts in inferred belief boundaries.
Across our experiments, LLMs maintain well-separated True and False representations, but both familiar and unfamiliar Neither statements occupy context-dependent regions. Unfamiliar Synthetic Neither statements induce the largest rotations and flip rates, showing that unfamiliar content disrupts an LLM’s veracity structure more than familiar Fictional content. Together, these analyses provide a systematic and principled way to evaluate the robustness of LLM veracity representations under different semantic assumptions, an essential step toward diagnosing and mitigating factual inconsistency.
@@
    Data: We introduce a new dataset of fictional statements across three factual domains, enabling controlled comparisons between familiar (Fictional) and unfamiliar (Synthetic) Neither content.
@@
    Method: We introduce and study representational stability in LLMs by combining activation-based probes with controlled label perturbations that vary the operational definition of truth.
@@
    Representational Structure: We show across sixteen open-source LLMs that True and False activations form tightly aligned clusters, while Neither statements (familiar and unfamiliar) occupy distinct regions, reflecting differences in training familiarity rather than superficial linguistic form.
  4. 4.
    Stability: We show that the unfamiliar Synthetic statements produce the largest rotations in the truth directions and the highest prediction flip rates (up to $40\%$ in Word Definitions), indicating that previously unseen yet semantically factual content most strongly destabilizes veracity geometry.
@@ 2 Related Work
Understanding how LLMs encode veracity touches on three research threads: (1) representation-based probing, (2) in-context stability, and (3) epistemic distinctions between belief, knowledge, and fact. We connect these strands by examining how familiar and unfamiliar Neither statements perturb the latent veracity geometry, thereby assessing how stable an LLM’s representations are under shifts in semantic assumptions.
Table 1: Summary of datasets and statement types. Number of affirmative (A) and negated (N) statements for each type across the three datasets, along with examples. Each dataset includes True, False, Synthetic, and Fictional statements, while Noise consists of randomly generated Gaussian activation vectors matched in dimensionality and distribution to the real statement embeddings. Synthetic statements serve as Neither statements that were not seen during LLM training, while Fictional statements are familiar Neither statements.
@@
(Fi) Snozzberry is a type of a berry.
A version of this table without the Fictional and Noise columns can be found in [savcisens2025trilemma].
Representation-based probing methods determine which properties are linearly recoverable from hidden states, revealing what models represent beyond input-output behavior [conneau2018you, hewitt2019structural, bert_probing]. Much of this work focuses on linguistic or syntactic recoverability, but recent studies have examined geometric structure, specifically, whether True and False statements form separable or directionally aligned clusters in activation space [marks2310geometry, burger2024truth]. Of particular relevance is the sAwMIL framework [savcisens2025trilemma], which uses multiple-instance learning and conformal prediction to classify statements as True, False, or Neither. Hallucination-detection studies likewise suggest that hidden states encode strong veracity signals even when outputs are incorrect [han-etal-2025-simple]. These empirical approaches complement philosophical work clarifying when neural components should count as representations [harding2023operationalizing].
A separate line of research demonstrates that LLM outputs are highly sensitive to prompting and context. Models are vulnerable to jailbreaks [wei2023jailbroken], sycophancy [sharma2024towards], word variation [elazar2021measuring], and multi-turn drift [li2025firm]. These studies diagnose behavioral brittleness rather than instability in the underlying representations, though recent evidence on epistemic familiarity suggests that some output-level failures may trace back to deeper weaknesses in internal epistemic structure [kadavath2022language].
LLMs also struggle to distinguish between belief, knowledge, and fact. Suzgun et al. [suzgun2025language] show that LLMs often fail to track agents’ beliefs when those beliefs are mistaken, underscoring weaknesses in their epistemic structure. Uncertainty-focused analyses reveal similar failures under epistemic ambiguity, i.e., when information admits multiple plausible interpretations [abbasi2024believe]. Theoretical work also argues that the study of LLM beliefs lacks unified standards, with Herrmann and Levinstein proposing criteria for when internal states should count as belief-like [herrmann2024standards]. Formal epistemology offers a complementary perspective: Leitgeb’s theory of $P$-stability links rational belief to stable truth assignments under small contextual changes [leitgeb2014stability]. This connection motivates our focus on representational stability as an epistemic property of model activations rather than model outputs.
We bridge probing and in-context work by measuring how controlled label perturbations reshape inferred truth directions, thereby identifying which types of familiar and unfamiliar Neither statements most strongly disrupt an LLM’s latent encoding of veracity.
@@ 3 Methodology
We treat an LLM’s internal activations as a proxy for its belief structure and use the decision boundary learned by a linear probe as a geometric diagnostic of how truth is encoded in that space. We define representational stability as the consistency of this boundary under controlled perturbations of what counts as True (see Fig. 1). The notation introduced in this section is summarized in Supplementary Table A1.
Table 2: Label configurations for experiments. Each row defines the composition of the True and Not True classes used when retraining the probe under different perturbation conditions. The baseline (Original) probe is trained on True statements versus all others, while perturbed probes redefine the True class to include additional statement types to test representational stability under shifts in the operational definition of truth. Fictional(T) denotes fictional truth and Fictional(F) denotes fictional falsehood.
Perturbation Type Statements Considered True Statements Considered Not True
@@
Original Baseline True False + Synthetic + Fictional + Noise
Perturbed Synthetic True + Synthetic False + Fictional + Noise
Fictional True + Fictional False + Synthetic + Noise
Fictional(T) True + Fictional(T) False + Synthetic + Fictional(F) + Noise
Noise True + Noise False + Synthetic + Fictional
Statements and Labels.
We begin with a collection of statements $\mathcal{S}=\{s_{i}\}_{i=1}^{N}$ drawn from factual domains. Each statement has a ground-truth veracity label $y_{i}\in\{\texttt{True},\texttt{False},\texttt{Neither}\}$. The Neither category includes familiar Fictional statements (e.g., “Gotham City is in New Jersey”) and unfamiliar Synthetic fact-like statements (e.g., “The city of Norminsk is located in Jamoates”).
Activation Extraction.
Each statement $s_{i}$ is passed through an LLM $\mathcal{M}$ to obtain its internal activation $\mathbf{z}_{i}^{(l)}$ a layer $l$, chosen empirically to maximize linear separability between True and Not True statements [marks2310geometry, burger2024truth, savcisens2025trilemma]. These activations constitute the model’s veracity representations. We construct the dataset
$\mathcal{D}=\{(\mathbf{z}_{i}^{(l)},y_{i})\}_{i=1}^{N},$
and partition it into training, calibration, and test splits: $\mathcal{D}_{\mathrm{train}}$, $\mathcal{D}_{\mathrm{cal}}$, and $\mathcal{D}_{\mathrm{test}}$.
Baseline Probe Training.
We train a max-margin, multiple-instance probe $\mathcal{P}$ (namely, sAwMIL [savcisens2025trilemma]) that learns a linear decision boundary $f(\mathbf{z})=\vec{w}\cdot\mathbf{z}+b$, where $\vec{w}$ represents the truth direction. Labels are encoded as $y^{\prime}_{i}\in\{+1,-1\}$:
$y^{\prime}_{i}=\begin{cases}+1,&\text{if }y_{i}=\texttt{True}\\ -1,&\text{if }y_{i}=\texttt{Not True}\end{cases}$
Training on $\mathcal{D}_{\mathrm{train}}$ yields the baseline classifier. The set of statements that the probe, and by extension the LLM’s latent geometry, represents as True is
$\mathcal{B}_{\mathrm{true}}=\{\,s_{i}\mid\mathbf{z}_{i}^{(l)}\in\mathcal{D}_{\mathrm{test}},\ y_{i}=\texttt{True},\ \hat{y}_{i}=+1\,\}.$
We interpret $\mathcal{B}_{\mathrm{true}}$ as the model’s baseline belief set under the original definition of truth.
Label Perturbations and Retraining.
To assess representational stability, we systematically vary which Neither statements are treated as True. Let $\mathcal{N}$ denote all Neither statements and partition it into two subsets $\mathcal{N}_{1}$ and $\mathcal{N}_{0}$, such that
$\mathcal{N}=\mathcal{N}_{1}\cup\mathcal{N}_{0},\qquad\mathcal{N}_{1}\cap\mathcal{N}_{0}=\emptyset.$
Relabeling $\mathcal{N}_{1}$ as True and $\mathcal{N}_{0}$ as Not True yields the perturbed labels
$y^{\prime\prime}_{i}=\begin{cases}+1,&\text{if }y_{i}\in\{\texttt{True}\}\cup\mathcal{N}_{1},\\ -1,&\text{if }y_{i}\in\{\texttt{False}\}\cup\mathcal{N}_{0}.\end{cases}$
Using the same data splits, we retrain the probe to obtain $(\vec{w}^{\prime},b^{\prime})$ and define the perturbed belief set as
$\mathcal{B}^{\prime}_{\mathrm{true}}=\{\,s_{i}\mid\mathbf{z}_{i}^{(l)}\in\mathcal{D}_{\mathrm{test}},\ y_{i}=\texttt{True},\ \hat{y}^{\prime}_{i}=+1\,\}.$
Comparing $\mathcal{B}_{\mathrm{true}}$ and $\mathcal{B}^{\prime}_{\mathrm{true}}$ reveals how the probe’s truth assignments shift under modified semantic assumptions while the underlying LLM activations remain fixed.
Quantifying Representational Stability.
Stability is quantified in two complementary ways. First, we measure geometric stability by comparing the decision boundaries $(\vec{w},b)$ and $(\vec{w}^{\prime},b^{\prime})$. Cosine similarity between $\vec{w}$ and $\vec{w}^{\prime}$ captures rotational changes in the truth direction, while $|b-b^{\prime}|$ captures translational shifts of the hyperplane.
Second, we evaluate prediction stability by comparing belief sets:
$\mathcal{B}_{\mathrm{true}}\cap\mathcal{B}^{\prime}_{\mathrm{true}}$ $\text{(stable truths)},$
$\mathcal{B}_{\mathrm{true}}\setminus\mathcal{B}^{\prime}_{\mathrm{true}}$ $\text{(epistemic retractions)},$
$\mathcal{B}^{\prime}_{\mathrm{true}}\setminus\mathcal{B}_{\mathrm{true}}$ $\text{(epistemic expansions)}.$
Following $P$-stability theory [leitgeb2014stability], retractions indicate stronger instability because they withdraw beliefs, whereas expansions reflect milder over-inclusiveness.
Together, these measures assess how reliably the LLM’s veracity representations support stable truth assignments under shifts in semantic boundaries. The probe’s decision boundary thus serves not as an end task, but as a geometric lens on how belief, falsity, and plausibility are structured within the LLM’s activation space.
Figure 2: Character bigram distributions of statements. Rank–frequency plots of normalized character bigram counts for True (green), False (red), Synthetic (yellow), and Fictional (blue) statements in the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. For each dataset, we compute per-type bigram frequencies, normalize within type, sort bigrams by their frequency under True statements, and plot log-normalized frequency with a moving-average smoothing. Across datasets, the True, False, and Synthetic distributions are nearly indistinguishable, whereas the Fictional distribution decays more slowly, marking it as structurally distinct.

### v3

# Representational and Behavioral Stability of Truth in Large Language Models

@@ Abstract
Large language models (LLMs) are increasingly used as information sources, yet small changes in semantic framing can destabilize their truth judgments. We propose P-StaT (Perturbation Stability of Truth), an evaluation framework for testing belief stability under controlled semantic perturbations in representational and behavioral settings via probing and zero-shot prompting. Across sixteen open-source LLMs and three domains, we compare perturbations involving epistemically familiar Neither statements drawn from well-known fictional contexts (Fictional) to those involving unfamiliar Neither statements not seen in training data (Synthetic). We find a consistent stability hierarchy: Synthetic content aligns closely with factual representations and induces the largest retractions of previously held beliefs, producing up to $32.7\%$ retractions in representational evaluations and up to $36.3\%$ in behavioral evaluations. By contrast, Fictional content is more representationally distinct and comparatively stable. Together, these results suggest that epistemic familiarity is a robust signal across instantiations of belief stability under semantic reframing, complementing accuracy-based factuality evaluation with a notion of epistemic robustness.
@@ 1 Introduction
Figure 1: Overview of the P-StaT framework. P-StaT evaluates the stability of an LLM $\mathcal{M}$’s beliefs by using a decision function $g$ to map statements $s$ to True or Not True under semantic assumptions $\Theta$. (a) In the baseline case $\Theta_{0}$, only True statements (purple) are labeled True, while False (green) and Neither (orange) statements are labeled Not True. Under a perturbed assumption set $\Theta$, a subset of Neither statements is treated as True. (b) The sets of True statements that $\mathcal{M}$ believes to be True in the baseline ($\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}$) and perturbed ($\mathcal{B}^{\Theta}_{\mathrm{true}}$) cases are identified via $g$. (c) Comparing these belief sets reveals unstable beliefs that are either retracted ($\mathcal{R}$) or expanded ($\mathcal{E}$). Crucially, $g$ can be instantiated using either representation-based probing or behavioral zero-shot prompting, enabling direct connections between instability in latent space and behavioral instability.
@@
Large language models (LLMs) are increasingly used as sources of information, but their behavior often blurs the line between knowledge and plausibility [2, 12]. While humans are expected to distinguish between True, False, and Neither-valued statements, it remains unclear whether LLMs form similarly structured internal representations, or whether such representations predict when truth judgments remain stable under changes in semantic framing [1, 26]. This gap is increasingly consequential as LLMs are deployed in high-stakes informational settings, where instability can undermine reliability and contribute to undesirable behaviors such as hallucinations [19, 16].
Existing work approaches this problem from two largely separate directions. Representation-based probing studies ask whether True and False statements occupy separable regions in activation space and whether “truth directions” transfer across LLMs or domains [20, 5, 23]. Behavioral studies document that small changes in prompting, context, or adversarial framing can substantially alter an LLM’s outputs, including its apparent confidence and self-consistency [29, 9, 18, 30]. While both lines of work reveal important failure modes, they are rarely connected through a shared experimental lens. Thus, we lack a unified framework for testing whether a candidate stability signal in latent space corresponds to stability in LLM behavior [13, 14].
We address this gap by proposing the Perturbation Stability of Truth framework, P-StaT, which links internal representations and LLM behavior through the lens of belief stability (Fig. 1). P-StaT asks how stable LLM truth judgments remain when semantic assumptions are systematically varied. This perspective is motivated by formal epistemology: under Leitgeb’s notion of $P$-stability, a belief system is epistemically well-formed only if established beliefs are preserved under small, principled changes in evidential context [17].
We operationalize semantic variation through perturbations $\Theta$ that modify which Neither statements are treated as compatible with truth. We distinguish epistemically familiar Neither statements (Fictional) from unfamiliar Neither statements (Synthetic) across three domains. Crucially, the same perturbation is instantiated both representationally, by retraining probes, and behaviorally, via zero-shot prompting.
We observe a consistent pattern: Unfamiliar Synthetic content induces the largest epistemic retractions in both representational and behavioral settings. By contrast, familiar Fictional content occupies more distinct regions of activation space and produces substantially fewer retractions. Together, these results identify epistemic familiarity as a cross-instantiation signal for when LLM truth judgments are most likely to destabilize under semantic reframing.
@@
    We introduce a dataset of fictional statements across three domains (City Locations, Medical Indications, and Word Definitions) to enable controlled comparisons between epistemically familiar and unfamiliar Neither content.
@@
    We propose P-StaT, a perturbation-based stability framework grounded in epistemic notions of belief stability that supports both probing and zero-shot prompting.
@@
    Across sixteen open-source LLMs, we show that unfamiliar Synthetic statements induce the largest epistemic retractions both representationally and behaviorally, identifying epistemic familiarity as a key signal of stability.
@@ 2 Related Work
Our work lies at the intersection of three research threads: (i) representation-based probing of veracity, (ii) behavioral sensitivity to semantic context, and (iii) epistemic notions of belief stability.
Representational Probing of Veracity.
Probing examines which properties are recoverable from hidden representations of LLMs, providing insights into what LLMs encode beyond observable behavior [7, 15, 27]. Recent work has shown that True and False statements form separable clusters across LLMs and domains in activation space [20, 5]. The sAwMIL framework [23] extends this work by modeling True, False, and Neither statements as distinct representational directions. Related work on hallucination detection similarly suggests that hidden representations encode strong veracity signals even when outputs are wrong [12].
Behavioral Stability and In-Context Effects.
A separate line of research documents small changes in wording or framing can lead to large shifts in LLM behavior, including jailbreak vulnerabilities [30], sycophancy [25], and instability under paraphrasing [9]. Recent work has framed this brittleness in terms of behavioral consistency across interactions. For example, Li et al. introduce benchmarks for evaluating response stability in multi-turn settings [18]. Related work on in-context learning suggests that larger LLMs can override pretraining-induced semantics when presented with conflicting in-context labels, while smaller LLMs rely more heavily on prior knowledge [31]. Complementarily, Bigelow et al. propose a Bayesian account in which both prompting and activation steering modify latent beliefs [4].
Epistemic Belief Stability.
LLMs also exhibit systematic difficulties in tracking epistemic distinctions such as belief, knowledge, and factuality [26, 1]. Herrmann and Levinstein argue that the study of LLM beliefs lacks unified criteria for when internal representations should count as belief-like [14]. Formal epistemology offers a complementary perspective: Leitgeb’s notion of $P$-stability characterizes rational belief systems as those that preserve established beliefs under small, justified changes in evidential context [17].
By combining representational probing with behavioral perturbations through P-StaT, our work bridges prior studies of latent veracity structure and in-context sensitivity. We assess how beliefs respond to principled shifts in semantic assumptions, providing a unified view of epistemic stability across representations and behavior.
@@
(Fi) Snozzberry is a type of berry.
Table 1: Summary of datasets and statement types. Number of affirmative (A) and negated (N) statements across the three datasets, along with examples. Each dataset includes True, False, Synthetic, and Fictional statements, while Noise consists of randomly generated Gaussian activation vectors matched in dimensionality and distribution to the statement embeddings. Synthetic statements serve as Neither statements that were not seen during LLM training, i.e., $\mathcal{N}_{\mathrm{unf}}$, while Fictional statements are familiar Neither statements $\mathcal{N}_{\mathrm{fam}}$. A version of this table without the Fictional and Noise columns can be found in [23].
@@ 3 Methodology
We study how LLMs organize statements in activation space and whether this organization predicts the stability of beliefs under semantic reframing using P-StaT (Perturbation Stability of Truth), which applies the same perturbations in both probing and zero-shot evaluations.
3.1 Epistemic Familiarity and Neither Statements
Let $\mathcal{S}=\{s_{i}\}_{i=1}^{N}$ denote declarative statements drawn from factual domains with labels $y_{i}\in\{\texttt{True},\texttt{False},\texttt{Neither}\}$. Let $\mathcal{N}=\{s_{i}\in\mathcal{S}\mid y_{i}=\texttt{Neither}\}$ denote the set of Neither statements. While all statements in $\mathcal{N}$ lack real-world truth value, $\mathcal{N}$ is heterogeneous and can be partitioned into familiar (Fictional) and unfamiliar (Synthetic) subsets:
$\mathcal{N}=\mathcal{N}_{\mathrm{fam}}\cup\mathcal{N}_{\mathrm{unf}},\qquad\mathcal{N}_{\mathrm{fam}}\cap\mathcal{N}_{\mathrm{unf}}=\emptyset.$
Here, $\mathcal{N}_{\mathrm{fam}}$ contains well-known fictional entities likely present in training corpora, while $\mathcal{N}_{\mathrm{unf}}$ contains constructed entities intended to be absent from training data. For Fictional, we additionally distinguish canonically true and false subsets, $\mathcal{N}_{\mathrm{fam}}^{(T)}$ and $\mathcal{N}_{\mathrm{fam}}^{(F)}$.
We evaluate three domains that differ in how sharply truth and falsehood are delineated (City Locations, Medical Indications, Word Definitions; Table 1) [23, 20, 5]. True, False, and Synthetic statements originate in [23], while we introduce the new Fictional dataset. [^1] The new fictional dataset can be accessed at https://huggingface.co/datasets/samanthadies/representational_stability. Details on data construction and validation are in Supplementary Section B.
3.2 Veracity Representations
[…]
3.3 P-StaT: Perturbation Stability of Truth
P-StaT evaluates stability for a fixed LLM $\mathcal{M}$ by assuming the evaluation procedure
$g:(s_{i},\Theta)\mapsto\{\texttt{True},\texttt{Not True}\},$
where $\Theta$ encodes semantic assumptions about which statements should be treated as compatible with truth. The $g(\cdot,\Theta)$ interface is the core abstraction of P-StaT, enabling the same semantic perturbation to be instantiated in distinct evaluation settings.
Representational vs. Behavioral Stability.
In representational experiments, $g$ is implemented as $g_{\mathrm{p}}(s,\Theta)=h_{\Theta}(\phi_{\mathcal{M},l}(s))$, where $h_{\Theta}$ is a linear probe trained with labels induced by $\Theta$. In behavioral experiments, $g$ is implemented as $g_{\mathrm{zs}}(s,\Theta)$ via zero-shot prompting and a belief context $C_{\Theta}$. Both settings support the same perturbations, enabling direct comparison.
Baseline evaluation.
Under a baseline semantic interpretation $\Theta_{0}$, $g$ identifies the set of ground-truth True statements that are assigned True:
$\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}\;=\;\{\,s_{i}\mid y_{i}=\texttt{True},\;g(s_{i},\Theta_{0})=\texttt{True}\,\}.$
Perturbations.
A perturbation modifies the semantic assumptions encoded in $\Theta$. Let $\mathcal{N}_{\Theta}\subseteq\mathcal{N}$ denote the subset of Neither statements treated as compatible with truth under a perturbed interpretation. The perturbed interpretation $\Theta$ differs from $\Theta_{0}$ only in this reassignment and yields the perturbed belief set:
$\mathcal{B}^{\Theta}_{\mathrm{true}}\;=\;\{\,s_{i}\mid y_{i}=\texttt{True},\;g(s_{i},\Theta)=\texttt{True}\,\}.$
Thus, only the semantic interpretation of Neither content is altered.
Epistemic retractions.
Within P-StaT, we quantify stability by comparing the baseline and perturbed belief sets. Epistemic retractions, $\mathcal{R}\;=\;\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}\setminus\mathcal{B}^{\Theta}_{\mathrm{true}}$, capture ground-truth True statements that lose belief status under perturbation. While we also consider epistemic expansions, $\mathcal{E}\;=\;\mathcal{B}^{\Theta}_{\mathrm{true}}\setminus\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}$, for completeness, epistemic retractions are a stronger signal of belief instability than expansions because they withdraw previously held beliefs [17].
Condition $\mathcal{N}_{\Theta}$ Probing: Training Labels Zero-shot: Belief Context $C_{\Theta}$
@@ 3.3 P-StaT: Perturbation Stability of Truth
Baseline $\emptyset$ True vs. False+Synthetic+Fictional+Noise None
Synthetic $\mathcal{N}_{\mathrm{unf}}$ True+Synthetic vs. False+Fictional+Noise Synthetic
Fictional $\mathcal{N}_{\mathrm{fam}}$ True+Fictional vs. False+Synthetic+Noise Fictional
Fictional (T) $\mathcal{N}_{\mathrm{fam}}^{(T)}$ True+Fictional(T) vs. False+Synthetic+Fictional(F)+Noise Fictional (T)
Noise N/A True+Noise vs. False+Synthetic+Fictional N/A
Table 2: Perturbation conditions $\Theta$ and instantiations in P-StaT. Each condition corresponds to a semantic interpretation $\Theta$ defined by $\mathcal{N}_{\Theta}$, the Neither statements treated as compatible with truth. The same $\Theta$ is instantiated (i) representationally by retraining a probe with labels induced by $\mathcal{N}_{\Theta}$ and (ii) behaviorally by constructing a belief context $C_{\Theta}$ from the corresponding training statements. Noise is probing-only since it is a non-semantic control.